# Generative Location Modeling for Spatially Aware Object Insertion

### v1

@@ Abstract
Generative models have become a powerful tool for image editing tasks, including object insertion. However, these methods often lack spatial awareness, generating objects with unrealistic locations and scales, or unintentionally altering the scene background. A key challenge lies in maintaining visual coherence, which requires both a geometrically suitable object location and a high-quality image edit. In this paper, we focus on the former, creating a location model dedicated to identifying realistic object locations. Specifically, we train an autoregressive model that generates bounding box coordinates, conditioned on the background image and the desired object class. This formulation allows to effectively handle sparse placement annotations and to incorporate implausible locations into a preference dataset by performing direct preference optimization. Our extensive experiments demonstrate that our generative location model, when paired with an inpainting method, substantially outperforms state-of-the-art instruction-tuned models and location modeling baselines in object insertion tasks, delivering accurate and visually coherent results.
@@ 1 Introduction
Explicit modeling of object locations has recently proven to be an effective strategy for generating complex scenes. Methods that use this strategy often separate the tasks of determining where objects should be placed and what those objects should look like (Feng et al., 2024a; Phung et al., 2024; Cho et al., 2024; Lian et al., 2023). This two-step approach typically involves generating a spatial layout and then conditioning an image generation model on this layout to create the final image. The success of these approaches has brought renewed attention to location modeling as a key component for achieving realistic and coherent scene generation.
We explore explicit location modeling for object insertion, which involves determining where new objects can be placed in a given scene. This is an important task with applications in generative data augmentation (Zhao et al., 2023; Kupyn & Rupprecht, 2024; Fang et al., 2024), virtual reality (Park et al., 2005) and robotics (Cheong et al., 2020). Given a scene image and an instruction specifying an object to be added, object insertion aims to place the object in the scene with a realistic appearance and geometry, while preserving the background and other objects. Unlike generating a full scene layout (Gupta et al., 2021), object insertion must account for the rich contextual information provided by the image, which imposes strong constraints on where objects can and cannot be placed.
Current state-of-the-art object insertion methods rely on instruction-tuned image editing (Brooks et al., 2023; Zhao et al., 2024; Zhang et al., 2024b; Wasserman et al., 2024), which requires models to jointly learn the realism in both appearance (what) and spatial placement (where). This is a difficult task that requires large-scale object insertion datasets that are inherently hard to construct. Despite training on such datasets, existing approaches often favor realism in appearance over placement, generating objects in unrealistic locations, replace or destroy existing objects, and create unintended changes in unrelated areas, as seen in Figure 1 (a).
In this work, we propose a two-stage object insertion pipeline that first identifies the object location using a dedicated location model, then generates the object locally in the location using an inpainting model, as illustrated in Figure 1 (b). Object placement datasets (Liu et al., 2021; Wasserman et al., 2024) suitable for training such models have recently been released. As it is unfeasible to manually label all possible locations, these datasets are inherently sparse, providing annotations for less than 1% of possible locations. Additionally, there may be multiple realistic locations for any given image, making location modeling a one-to-many problem. State-of-the-art object placement models handle these issues by either assuming unlabeled areas as implausible locations (Lin et al., 2018; Tripathi et al., 2019; Zhang et al., 2020) or crafting custom loss functions (Niu et al., 2022; Zhu et al., 2023). However, these strategies risk penalizing unlabeled positive locations and are also highly sensitive to annotation sparsity, limiting their scalability across datasets.
To overcome these challenges, we approach the problem from a generative perspective, as training a generative model only requires samples from the target distribution. Specifically, we represent the input image and the object class as a sequence of tokens and use an autoregressive transformer model (Vaswani et al., 2017; Radford et al., 2019) to iteratively decode bounding box coordinates of plausible object locations. Furthermore, we can fully leverage any additional negatively labeled locations by treating pairs of positive and negative labels as a preference dataset, where positive locations are preferred over negative locations. This strategy enables to perform direct preference optimization (Rafailov et al., 2024) on the location model, further enhancing its accuracy.
We empirically observe that given a precise location of the object to be generated, off-the-shelf inpainting models outperform state-of-the-art instruction-tuned object insertion models. Moreover, we observe that any inaccuracy in estimating the object location can significantly degrade the image quality, highlighting the role of location modeling as the critical component of the object insertion pipeline. Our location model produces high-quality locations, allowing it to outperform both instruction-finetuned models in object insertion and existing location models in terms of positional accuracy. We further validate the effectiveness of our approach in a user study.
Figure 1: Our proposed pipeline for object insertion (b), in contrast to instruction-tuned methods (a). We use a pretrained inpainting model and provide it with plausible locations for object insertion.
Summarizing, our main contributions are as follows:
  • We propose a two-stage object insertion approach, and overcome the limitations of instruction-tuned editing with an explicit location model.
  • We effectively handle sparsity in location annotations with a generative approach: an image-conditioned autoregressive transformer that models bounding box locations.
  • We can leverage available negative annotations following our generative formulation using direct preference optimization, further improving the accuracy of our location model.
  • Through extensive experiments and a user study, we demonstrate that our approach achieves state-of-the-art performance in location modeling, and significantly outperforms instruction-tuned image editing models in object insertion tasks.
@@ 2 Related Work
Instruction-based Image Editing. InstructPix2Pix (Brooks et al., 2023) introduced an approach that finetunes Stable Diffusion (Rombach et al., 2021) to interpret text instructions, trained with a dataset of paired images before and after specific edits. Subsequently, other methods (Zhang et al., 2024b; Hui et al., 2024; Zhang et al., 2024a) have released similar datasets for instruction-tuning, capable of changing the style and content of a given image.
Only recently has there been an emphasis on adding objects using text instructions, supported by datasets that provide paired images where specific objects have been artificially removed by inpainting (Wasserman et al., 2024; Zhao et al., 2024). One downside of these datasets is that they contain inpainting artifacts, which can cause models trained on them to replace existing objects or alter backgrounds. Additionally, since these models are typically trained to regenerate the entire image, they often introduce unintended changes to the scene. In contrast, our approach decouples the object insertion process by using a dedicated location model for determining placement and an inpainting model for rendering. This factorization allows for more control and precision in object insertion.
Object Placement. Traditionally, object placement has relied on copy-pasting an object segment by simply determining its location and scale (Zhang et al., 2020; Zhu et al., 2023; Zhou et al., 2022; Tripathi et al., 2019; Niu et al., 2022). However, this approach is not ideal for inserting objects into background images, as it requires the user to prepare an image of the object that fits seamlessly into the scene purely by placement.
@@
To avoid relying on object segments, some approaches predict locations from class labels by querying a classifier (Dvornik et al., 2018) on random bounding boxes, categorizing them as either plausible or implausible locations for the given class. To enable training a discriminative model, unlabeled locations are typically treated as negatives or implausible locations. Such an assumption is not always valid in object placement, as the lack of annotations for specific locations does not necessarily indicate they are implausible. Therefore, penalizing them may result in an inaccurate location model. To overcome this limitation, we instead propose a generative approach that requires only samples from the target distribution, and does not make any assumption about unlabeled locations and only requires samples of the target distribution (i.e., positive locations).
@@
Layout Generation. Another related category of work is focused on scene layout generation, usually via a generative model of bounding box locations (Jyothi et al., 2019; Gupta et al., 2021; Chai et al., 2023; Inoue et al., 2023) or segmentation maps (Lee et al., 2018). Such layouts are often used as a condition for image generation models such as GLIGEN (Li et al., 2023) to generate complex scenes (Feng et al., 2024a; Phung et al., 2024; Cho et al., 2024; Lian et al., 2023; Gani et al., 2024; Feng et al., 2024b). Unlike these full-layout generation approaches, our goal is to insert an object into an existing scene. Having access to the background image fundamentally changes the nature of the task. On one hand, the background image provides valuable context for realistic object placement, but on the other hand, it imposes strong constraints on where objects can be placed. We therefore design a location model that integrates these contextual cues and avoids placement in unrealistic locations.
@@ 3.1 Generative Location Modeling
To model this generative process, we train an autoregressive model (Vaswani et al., 2017) that sequentially predicts the bounding box coordinates of plausible locations. Specifically, each location $Y_{i}$ is represented as a bounding box with four components $\left[b_{1}^{i},\,b_{2}^{i},\,b_{3}^{i},\,b_{4}^{i}\right]=\left[x_{1},\,y_{1},\,x_{2},\,y_{2}\right]$, representing the coordinates of the top-left and bottom-right corners. Thus, given a dataset $\mathcal{D}$ which provides pairs of images $X$ and plausible locations $Y$ for object category $C$, we train the model using a negative log-likelihood objective:
@@
The model architecture and the training procedure are illustrated in Figure 2. We tokenize images and the class embeddings using a pre-trained Vision Transformer (ViT) (Dosovitskiy, 2021) and a CLIP encoder (Radford et al., 2021). We encode bounding box coordinates by quantizing them to a grid with equally spaced bins of 1 pixel wide, (i.e., $512$ location tokens for $512\times 512$ images). The image tokens and the target class token are prepended to the sequence, and our location model is trained to predict the probability of each coordinate in an autoregressive manner.
Figure 2: Training scheme during pretraining (left) and direct preference optimization (right).
3.2 Leveraging Negative Labels via Direct Preference Optimization
@@
While the training objective in Equation 2 allows training the location model on sparse positive annotations, training solely on positive feedback can lead to predictions in implausible locations. Incorporating negative annotations, when available, into the training objective can be beneficial for refining the model, encouraging it to assign lower likelihoods to undesirable locations and thereby improving overall accuracy.
@@
Our generative formulation allows us to use any negative labels in the dataset as well, so that the model can learn to avoid predicting bounding boxes for implausible locations. Specifically, we treat the positive and negative labels as a preference dataset, where positive locations are implicitly preferred over negative ones, even though annotators were not explicitly asked to rank them. Using this preference structure, we fine-tune the model with direct preference optimization (DPO) (Rafailov et al., 2024), penalizing high logits assigned to negative labels.
@@
Given a target location model $\pi_{\theta}$, a reference location model $\pi_{\text{ref}}$, and a preference dataset $\mathcal{D}_{\text{DPO}}$ where locations $Y^{+}$ are preferred over $Y^{-}$, we can derive a preference model (Bradley & Terry, 1952) as follows:
@@
$P(Y^{+}\succ Y^{-}\mid X,C)=\left(1+\exp\left(\beta\log\frac{\pi_{\theta}(Y^{-}\mid X,C)}{\pi_{\text{ref}}(Y^{-}\mid X,C)}-\beta\log\frac{\pi_{\theta}(Y^{+}\mid X,C)}{\pi_{\text{ref}}(Y^{+}\mid X,C)}\right)\right)^{-1}.$ (3)
@@
Here, $\pi_{\theta}$ and $\pi_{\text{ref}}$ output the logits for the target model and the reference model, respectively, and $\beta$ is a hyperparameter. We can maximize the preference of $Y^{+}$ locations by initializing the target and reference models from a pre-trained location model (trained using the objective in equation 2), and then optimizing the target model using a negative log-likelihood objective:
@@
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(Y^{\text{+}},Y^{\text{-}},X,C)\sim\mathcal{D}_{\text{DPO}}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(Y^{+}\mid X,C)}{\pi_{\text{ref}}(Y^{+}\mid X,C)}-\beta\log\frac{\pi_{\theta}(Y^{-}\mid X,C)}{\pi_{\text{ref}}(Y^{-}\mid X,C)}\right)\right].$ (4)
@@ 4 Experiments
Figure 3: Annotation format for the PIPE dataset (left) and OPA dataset (right). The PIPE dataset has one groundtruth location per image, whereas OPA provides multiple positive and negative locations.
@@ 4.1 Datasets and architecture
PIPE Dataset. The PIPE dataset (Wasserman et al., 2024) was created by removing objects from object detection datasets (Lin et al., 2014; Kuznetsova et al., 2020; Gupta et al., 2019) by inpainting. This process results in pairs of images, one including the object and the other without it. To train our location model we need positive bounding box locations, that we derive for the missing object by thresholding the pixel-wise difference between the two images. An example is reported in Figure 3 (left). While the dataset offers a large number of 888,000 samples, many images contain inpainting artifacts, potentially introducing noise in the bounding box extraction process.
@@
OPA Dataset. The OPA dataset (Liu et al., 2021) was created by asking human annotators to judge the plausibility of object placement locations for a subset of COCO images (Lin et al., 2014). This dataset includes on average 41.5 annotations per image, and can be used to encourage diversity in location model predictions (see Figure 3 right for an example). The train set includes 1022 images, and the test set include 130 images. As OPA provides negative labels, we also use OPA as the preference dataset for DPO training. Note that instruction-tuned editing models cannot leverage this location dataset, as they require fully rendered images for training.
@@
Implementation details. We use a small GPT-2 (Radford et al., 2019) architecture as our autoregressive location model. For tokenizing the image and the object class, we use a ViT (Dosovitskiy, 2021) model pre-trained on ImageNet-21K (Ridnik et al., 2021) as our image encoder and the CLIP text encoder (Radford et al., 2021) as our object class encoder. We pre-train on PIPE for 30K iterations and finetune on OPA for 3K iterations. For batch size 64, the model can be trained on a single Nvidia V100 GPU. We quantize each box coordinate into one of 512 bins (i.e., one bin per pixel). For DPO training, we train using the OPA dataset for 4K iterations. Please refer to Section D of the Appendix for further details.
@@
We evaluate the performance of our generative location model in an object insertion pipeline in Section 4.2, where we rely on PowerPaint (Zhuang et al., 2023) as the inpainting approach. We remark that our pipeline has the flexibility to incorporate any inpainting method for performing localized edits, and will ablate this choice in Section 4.4.

### v2

# Imagining the Unseen: Generative Location Modeling for Object Placement

@@ Abstract
Location modeling, or determining where non-existing objects could feasibly appear in a scene, has the potential to benefit numerous computer vision tasks, from automatic object insertion to scene creation in virtual reality. Yet, this capability remains largely unexplored to date. In this paper, we develop a generative location model that, given an object class and an image, learns to predict plausible bounding boxes for such an object. Our approach first tokenizes the image and target object class, then decodes bounding box coordinates through an autoregressive transformer. This formulation effectively addresses two core challenges in locatio modeling: the inherent one-to-many nature of plausible locations, and the sparsity of existing location modeling datasets, where fewer than 1% of valid placements are labeled. Furthermore, we incorporate Direct Preference Optimization to leverage negative labels, refining the spatial predictions. Empirical evaluations reveal that our generative location model achieves superior placement accuracy on the OPA dataset as compared to discriminative baselines and image composition approaches. We further test our model in the context of object insertion, where it proposes locations for an off-the-shelf inpainting model to render objects. In this respect, our proposal exhibits improved visual coherence relative to state-of-the-art instruction-tuned editing methods, demonstrating a high-performing location model’s utility in a downstream application.
Work completed during an internship at Qualcomm Technologies, Inc.Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
Identifying plausible locations for objects that do not exist in an image is an underexplored problem in computer vision. Traditional perception methods, such as object detectors [54, 36], primarily focus on detecting and localizing existing objects within scenes. However, numerous emerging applications demand a complementary capability, which is to determine where non-existing objects could realistically appear. For example, object insertion [55, 24, 45], generative data augmentation [50, 21, 8], content planning in virtual reality [31], and robotics scenarios [4] require models to reason about objects and their spatial relations, even if some of them are not yet present in the scene.
Figure 1: A location model identifies plausible “empty spaces” for a hypothetical object, given the name of the object. Images taken from the OPA dataset [^1] OPA images are under the MIT License: https://github.com/bcmi/Object-Placement-Assessment-Dataset-OPA/blob/main/LICENSE.
One can make this task explicit by training a location model. Training such a model poses significant challenges, as the task is inherently ambiguous and suffers from data scarcity. Unlike object detection, where existing objects are precisely annotated, annotations with plausible insertion locations typically lack exhaustive labels. Even, specialized datasets explicitly created for object placement [29, 43] annotate fewer than 1% of possible insertion points, as it is impractical to exhaustively label every region in the image. Previous approaches have often addressed this by either treating unlabeled regions as implicitly invalid [26, 40, 47] or relying on custom loss functions that heavily penalize unlabeled but potentially valid placements [30, 53]. Such methods are often sensitive to specific annotation distributions within datasets (e.g., number of annotation per image), hindering their generalization to newly released datasets.
In this paper, we overcome these limitations by recasting the problem from a generative perspective. Generative models naturally address data sparsity by directly modeling the distribution of plausible placements, enabling effective training from sparse positive examples without requiring exhaustive annotations. Specifically, we encode the input image along with the intended object class into a sequence of tokens, and train an autoregressive transformer [41, 33] to iteratively generate plausible bounding-box coordinates. Additionally, when negative annotations (indicating unsuitable locations) are available, we incorporate these as explicit preferences, favoring positive placements over negative ones through Direct Preference Optimization [35], which further improves the model’s accuracy.
Our experiments demonstrate that our location model consistently predicts more accurate and plausible object locations compared to existing discriminative methods, highlighting the advantages of a generative formulation. Some exemplar locations are illustrated in Footnote 1. To further showcase the practical utility of our model, we conduct comprehensive experiments in an object insertion scenario, where the predicted locations guide an off-the-shelf inpainting model to render new objects into existing scenes. We find that this approach outperforms recent image editing model for object insertion, highlighting the broader impacts of location models, which can potentially benefit downstream tasks, such as automated content creation, synthetic dataset construction, and planning for robotics.
In summary, our main contributions are:
  • We propose a generative transformer model that effectively identifies plausible object placement locations, naturally handling data sparsity and multimodal outcomes.
  • We leverage direct preference optimization as a method to incorporate negative labels, improving bounding-box prediction accuracy in when negative labels are available.
  • We demonstrate through extensive experiments that our method significantly surpasses existing techniques in terms of accuracy and realism, even in challenging object insertion scenarios.
@@ 2 Related Work
Figure 2: Training scheme of our generative location model during pretraining (left) and direct preference optimization (right).
Object Placement.
Traditionally, object placement has relied on copy-pasting an object segment by simply determining its location and scale [47, 53, 52, 40, 30]. However, this approach is limited because it assumes the object appearance and shape are already defined, which is often not the case when integrating new objects into diverse and dynamic environments.
@@
To avoid relying on object segments, some approaches predict locations from class labels by querying a classifier [7] on random bounding boxes, categorizing them as either plausible or implausible locations for the given class. To enable training a discriminative model (e.g., classifiers or object detectors), unlabeled locations are typically treated as negatives or implausible locations. Such an assumption is not always valid in object placement, as the lack of annotations for specific locations does not necessarily indicate they are implausible. Therefore, penalizing them may result in an inaccurate location model. To overcome this limitation, we instead propose a generative approach that requires only samples from the target distribution, and does not make any assumption about unlabeled locations and only requires samples of the target distribution (i.e., positive locations).
Layout Generation.
@@
Another related category of work is focused on scene layout generation, usually via a generative model of bounding box locations [18, 13, 3, 16] or segmentation maps [23]. Such layouts are often used as a condition for image generation models such as GLIGEN [24] to generate complex scenes [9, 32, 5, 25, 11, 10]. Unlike these full-layout generation approaches, our goal is to insert an object into an existing scene. Having access to the background image fundamentally changes the nature of the task. On one hand, the background image provides valuable context for realistic object placement, but on the other hand, it imposes strong constraints on where objects can be placed. We therefore design a location model that integrates these contextual cues and avoids placement in unrealistic locations.
Instruction-based Image Editing.
Some works choose to solve reasoning about object interactions implicitly, in context of another task, for example image editing. InstructPix2Pix [2] introduced an approach that fine-tunes Stable Diffusion [38] to interpret text instructions, trained with a dataset of paired images before and after specific edits. Subsequently, other methods [49, 14, 46] have released similar datasets for instruction-tuning, capable of changing the style and content of a given image.
Only recently has there been an emphasis on adding objects using text instructions, supported by datasets that provide paired images where specific objects have been artificially removed by inpainting [43, 51] or using training-free techniques [39]. One downside of training on such a dataset is that the data contains inpainting artifacts, which can cause models trained on them to replace existing objects or alter backgrounds. Additionally, since these models are typically trained to regenerate the entire image, they often introduce unintended changes to the scene. In contrast, a reliable location model enables us to decouple the object insertion process by first determining placement using the location model and render the object in that location using a inpainting model. This factorization allows for more control and precision in object insertion.
@@ 3.1 Generative Location Modeling
To model this generative process, we train an autoregressive model [41] that sequentially predicts the bounding box coordinates of plausible locations. Specifically, each location $Y_{i}$ is represented as a bounding box with four components $\left[b_{1}^{i},\,b_{2}^{i},\,b_{3}^{i},\,b_{4}^{i}\right]=\left[x_{1},\,y_{1},\,x_{2},\,y_{2}\right]$, representing the coordinates of the top-left and bottom-right corners. Thus, given a dataset $\mathcal{D}$ which provides pairs of images $X$ and plausible locations $Y$ for object category $C$, we train the model using a negative log-likelihood objective:
@@
The model architecture and the training procedure are illustrated in Fig. 2. We tokenize images and the class embeddings using a pre-trained Vision Transformer (ViT) [6] and a CLIP encoder [34]. We encode bounding box coordinates by quantizing them to a grid with equally spaced bins of 1 pixel wide, (i.e., $512$ location tokens for $512\times 512$ images). The image tokens and the target class token are prepended to the sequence, and our location model is trained to predict the probability of each coordinate in an autoregressive manner.
3.2 Leveraging Negative Labels via DPO
@@
While the training objective in Equation 2 allows training the location model on sparse positive annotations, training solely on positive feedback can lead to predictions in implausible locations. Incorporating negative annotations, whenever available, into the training objective can be beneficial for refining the model, encouraging it to assign lower likelihoods to undesirable locations and thereby improving overall accuracy.
@@
Our generative formulation allows us to use any negative labels in the dataset as well, so that the model can learn to avoid predicting bounding boxes for implausible locations. Specifically, we treat the positive and negative labels as a preference dataset, where positive locations are implicitly preferred over negative ones, even though annotators were not explicitly asked to rank them. Using this preference structure, we fine-tune the model with direct preference optimization (DPO) [35], penalizing high logits assigned to negative labels. We repeat the training objective below and refer the reader to Rafailov et al. [35] Eq. 1-6 for the full derivation.
@@
Given a target location model $\pi_{\theta}$ (i.e. the model currently being finetuned by DPO), a reference location model $\pi_{\text{ref}}$ (i.e. a frozen model trained by Equation 2 only), and a preference dataset $\mathcal{D}_{\text{DPO}}$ where locations $Y^{+}$ are preferred over $Y^{-}$, we can derive the likelihood of $Y^{+}$ being preferred over $Y^{-}$, based on how well the target location model $\pi_{\theta}$ predicts each location relative to $\pi_{\text{ref}}$, following the Bradley-Terry model [1]:
@@
$P(Y^{+}\succ Y^{-}\mid X,C)=\Biggl[1+\exp\Biggl(\\ \beta\left(\log\frac{\pi_{\theta}(Y^{-}\mid X,C)}{\pi_{\text{ref}}(Y^{-}\mid X,C)}-\log\frac{\pi_{\theta}(Y^{+}\mid X,C)}{\pi_{\text{ref}}(Y^{+}\mid X,C)}\right)\Biggr)\Biggr]^{-1}.$ (3)
@@
Here, $\pi_{\theta}$ and $\pi_{\text{ref}}$ output the logits for the target model and the reference model, respectively, and $\beta$ is a hyperparameter. We can maximize the preference of $Y^{+}$ locations by initializing the target and reference models from a pre-trained location model, and then optimizing the target model using a negative log-likelihood objective:
@@
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(Y^{+},Y^{-},X,C)\sim\mathcal{D}_{\text{DPO}}}\Biggl[\log\sigma\Biggl(\\ \beta\log\frac{\pi_{\theta}(Y^{+}\mid X,C)}{\pi_{\text{ref}}(Y^{+}\mid X,C)}-\,\beta\log\frac{\pi_{\theta}(Y^{-}\mid X,C)}{\pi_{\text{ref}}(Y^{-}\mid X,C)}\Biggr)\Biggr].$ (4)
@@ 4 Experiments
Figure 3: Annotation format for the PIPE [^2] PIPE images are under $c$-by-4.0 license: https://huggingface.co/datasets/choosealicense/licenses/blob/main/markdown/$c$-by-4.0.mdand OPA dataset. The PIPE dataset has one groundtruth location per image, whereas OPA provides multiple positive and negative locations.
@@ 4.1 Datasets and architecture
The PIPE dataset [43] was created by removing objects from object detection datasets [27, 22, 12] by inpainting. This process results in pairs of images, one including the object and the other without it. To train our location model we need positive bounding box locations, that we derive for the missing object by thresholding the pixel-wise difference between the two images. An example is reported in Footnote 2 (top). While the dataset offers a large number of 888,000 samples, many images contain inpainting artifacts, potentially introducing noise in the bounding box extraction process.
OPA Dataset.
@@
The OPA dataset [29] was created by asking human annotators to judge the plausibility of object placement locations for a subset of COCO images [27]. This dataset includes on average 41.5 annotations per image, and can be used to encourage diversity in location model predictions (see Footnote 2 (bottom) for an example). The train set includes 1022 images, and the test set include 130 images. As OPA provides negative labels, we also use OPA as the preference dataset for DPO training.
Implementation details.
@@
We use a small GPT-2 [33] architecture as our autoregressive location model. For tokenizing the image and the object class, we use a ViT [6] model pre-trained on ImageNet-21K [37] as our image encoder and the CLIP text encoder [34] as our object class encoder. We pre-train on PIPE for 30K iterations and fin-etune on OPA for 3K iterations. For batch size 64, the model can be trained on a single Nvidia V100 GPU. We quantize each box coordinate into one of 512 bins (i.e., one bin per pixel). For DPO training, we train using the OPA dataset for 4K iterations. Please refer to Section D of the Appendix for further details.
@@
We evaluate the performance of our generative location model in an object insertion pipeline in Section 4.3, where we rely on PowerPaint [55] as the inpainting approach. Specifically, we utilize the V2 version [^3] https://github.com/open-mmlab/PowerPaint, built on top of BrushNet [17]. We remark that our pipeline has the flexibility to incorporate any inpainting method for performing localized edits (see later in Section 4.3).
Figure 4: Example evaluation scenario. Predicted boxes are counted only if a positive or negative ground-truth box meets an IoU above the threshold.
4.2 Location Modeling
Baselines.
We compare our generative location model against discriminative approaches that classify locations as plausible or implausible. First, we train ContextCNN [7], a classifier designed to assess masked regions of an image, determining their suitability for object placement. To be able to use it with the OPA class space, we retrain the model on COCO (whose classes comprehend all the ones in OPA). Furthermore, we train a Faster-RCNN object detector [36] using positive OPA annotations to serve as a high-performing discriminative model.
Additionally, we compare to two object placement baselines [52, 30] that perform placement of specific object segments rather than generic class labels. We use the official implementations relying on foreground segments available within the OPA dataset. Since these methods are provided with the ground-truth object image and its aspect ratio, they are not a perfect baseline for our task. Nonetheless, we include their results for completeness as they also operate on the OPA dataset.
Evaluation Metrics.
The OPA test set provides plausible (positive) and implausible (negative) locations for objects given an image. However, due to the sparse nature of these annotations, it is impossible to sample locations until every ground-truth bounding box is matched with a prediction. Therefore, we evaluate location models based on a “hit rate” metric, which compares the rate of each predicted box being a plausible or implausible location. Specifically, we measure the True Positive Rate (TPR) and False Positive Rate (FPR) for a given set of predicted locations. Given $K$ predicted locations, we match them to the ground-truth labels using the Hungarian algorithm [20], where the cost function is the inverse of the intersection over union (IoU).
True positive predictions are defined as predictions assigned to positive labels with an IoU above 0.7, while false positive predictions are those assigned to negative labels under the same IoU threshold. Any positive or negative ground-truth locations that are not matched are counted as false negatives and true negatives. Predictions that do not correspond to any labeled locations are ignored, as their true labels cannot be determined. We refer the reader to Figure 5 for an example of such assignments. TPR and FPR are then computed using standard definitions:
$\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}},\quad\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}.$ (5)
Intuitively, TPR represents the rate of a predicted location being a correct (positive) location, and FPR represents the rate of a predicted location being an incorrect (negative) location.
Figure 5: TPR-FPR curves. Each line is constructed by sampling $\{10,20,\ldots,100\}$ locations. Top-left is better.
Results.
We plot the TPR and FPR for different number of sampled locations $K=10,20,\dots,100$ in Figure 5. Our generative location model consistently achieves a higher TPR at the same FPR, appearing in the top-left region of the plot. In contrast, discriminative baselines, ContextCNN [7] and FasterRCNN [36] fail to reach a high TPR.. This is likely due to the penalization of unlabeled positive locations during training, which limits their ability to generalize effectively. Our generative approach, on the other hand, avoids assumptions about unlabeled locations, enabling it to identify plausible placements with greater accuracy. This result highlights the effectiveness of generative modeling, particularly in scenarios where annotation sparsity hinders the training of discriminative models.
We additionally compare two object placement methods, FOPA [30] and GracoNet [52]. Unlike our setting, these approaches do not predict bounding boxes given object classes (which have four degrees of freedom: the coordinates of the bounding box), but rather predict the location and scale of an object (with three degrees of freedom) given an image of the object. Despite having access to the ground-truth object segments and aspect ratios, both FOPA and GracoNet fail to outperform our approach, further demonstrating the advantages of our generative model in producing more accurate and plausible placements.
Figure 6: Quantitative evaluation on the OPA dataset, higher and to the right is best. For instruction-tuned approaches, each dot represents a different guidance scale ranging from 2 to 10. For other methods, guidance scale has a negligible effect, hence we show a single point.